Consider a binomial experiment with n = 8 trials where the probability of success on a single trial is p = 0.15. (For each answer, enter a number. Round your answers to three decimal places.)

Find P(r = 0).
Answer #1

n = 8, p = 0.15, q = 1 - p = 0.85

P(r) = C(n, r) p^r q^(n - r)

P(0) = C(8, 0) 0.15^0 0.85^(8 - 0) = 0.272

Answer: 0.272

[Please give me a Thumbs Up if you are satisfied with my answer. If you are not, please comment on it, so I can edit the answer. Thanks.]